# Rate, Explain and Cite (REC): Enhanced Explanation and Attribution in Automatic Evaluation by Large Language Models

### v2

@@ 6.3 General Capabilities
The RewardBench dataset assesses LLMs’ performance across several critical abilities using curated chosen-rejected response pairs Lambert et al. (2024). The evaluation involves determining the model’s win percentage based on its ability to score the “chosen” response higher than the “rejected” one. Specifically, it’s focused on testing chat performance, safety measures, and reasoning (both in terms of code and math skills), while controlling for potential biases such as overfitting to prior datasets.
We first create prompts following the RewardBench evaluation format, where each prompt consisting of a query and two potential responses. The task for LLMs is to evaluate the two responses and determine which is better. The performance of the reward models is quantified by calculating the win percentage for chosen-rejected pairs associated with each prompt. A “win” is defined by a scenario where an LLM presents a preference for the selected answer. This quantitative measure provides a clear benchmark for evaluating model performance across tasks. By averaging the results across the 4 core sections: Chat, Chat Hard, Safety, and Reasoning, a comprehensive measure of each model’s win percentage is derived. As shown in Table 4, our REC models are able to achieve SOTA performance among all generative LLMs on the RewardBench Leaderboard, demonstrating their outstanding ability to serve as an autoevaluator for pairwise comparisons.
@@
This is a benchmark for measuring the grounding capabilities of autoevaluators. Given a reference document and a claim, the autoevaluator determines if the claim is fully supported by the document. This holistic benchmark combines 10 attribution datasets used in recent studies on LLM factuality. Given some of the models such as GPT-3.5 have smaller context windows, we removed examples longer than 16K tokens from our evaluation. We use the same prompt instructions across categories, as shown in Appendix G. To reduce model API costs, we randomly sampled 256 examples per evaluation task similar to the approach used in the FLAMe Vu et al. (2024b).
Table 5 presents our attribution results on LLM-AggreFact Tang et al. (2024b), categorized into four common use-cases: (1) LLM-FactVerify: fact verification of LLM-generated responses, (2) Wiki-FactVerify: evaluating correctness of Wikipedia claims, (3) Summarization: assessing faithfulness of summaries, and (4) Long-form QA: evaluating long-form answers to questions. REC-12B and REC-70B outperform all other models in all four categories, and REC-8B outperform all the other similar-sized models (Mistral-7B and Mistral-Nemo), demonstrating their strong performance in attribution on diverse datasets.
@@ 6.4 Bias Testing
Recent studies have found that LLM-as-a-Judge often exhibits cognitive biases, such as preferences for verbosity, egocentrism, bandwagon, and an overly authoritative tone Wang et al. (2024a); Koo et al. (2024); Zheng et al. (2024); Chen et al. (2024). To investigate the biases of the compared models, we evaluate them on the CoBBLEr benchmark (Cognitive Bias Benchmark for LLMs as EvaluatoRs) Koo et al. (2024). This dataset is designed to evaluate the quality and reliability of LLMs when used as automated evaluators in a question-answering (QA) setting. It assesses the presence of six cognitive biases, both implicit and induced, when LLMs are tasked with ranking responses generated by various other models. CoBBLEr’s core objective is to identify the extent of bias in LLM evaluation outputs.
The CoBBLEr dataset includes 50 QA instructions, randomly selected from two well-established benchmarks: BIG-bench bench authors (2023) and ELI5 Fan et al. (2019). 16 LLMs, both open and closed-source models, generate responses to these instructions. The evaluations involve pairwise comparisons between the responses of two models, wherein each model also acts as an evaluator to rank its own and others’ outputs. The biases tested are categorized into two groups: (1) Implicit biases, such as egocentric bias (where a model tends to prefer its own outputs), and (2) Induced biases, such as order bias, where the ranking of responses is influenced by their order in the evaluation.
@@ 7 Conclusions
LLMs excel in generating coherent and high-quality text for various tasks, but evaluating content quality is challenging due to issues like factual inaccuracies and hallucination. This paper introduces three fine-tuned models—REC-8B, REC-12B, and REC-70B—designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, completeness. These models not only provide ratings for these metrics but also offer detailed explanations and verifiable citation, i.e., Ratings, Explanations, and citation (REC), thereby enhancing trust in the content. To achieve this goal, a curated dataset for explanations and citation fine-tuning is proposed to facilitate future research, which is the first public dataset containing both content quality citation and RAG citation. The models support various citation modes, including Post-fix citation, Post-fix citation mode with snippet, Inline citation, and Inline citation mode with snippet to balance the trade-off between latency and granularity of citation. Extensive evaluations on diverse benchmarks—ALCE, ExpertQA, ABCD summarization, RewardBench, LLM-AggreFact, and CoBBLEr—demonstrate that the fine-tuned LLM autoevaluator, REC-70B, outperforms state-of-the-art models. It excels in content evaluation, delivering clear explanations and ensuring citation accuracy.
@@
Our work introduces the innovative Rate, Explain, and Cite (REC) model for trusted automated AI evaluations, offering significant benefits but with a few limitations that future research will address. First, our work is limited to the baselines we compare, as non-LLM citation methods, for instance, approaches that leverage sentence embedding Reimers and Gurevych (2019) similarity to identify and cite the most semantically relevant sources were not included. Including such techniques could serve as valuable baselines to further validate our model’s generalization and effectiveness. Second, the need for post-generation explanations and citation may introduce latency. Future work will compare latencies across models and explore optimizations to improve speed. We are also considering fine-tuning the snippet output using starting and ending character numbers to optimize for latency. Finally, even though the 12B base model is capable of generating multi-lingual output, our current evaluations are limited to English datasets. Expanding to multilingual evaluations is a key goal for future research.
Ethical Considerations and Risks
Our contributions present a novel general-purpose autoevaluator that supports various modes of citation, as well as a dataset with citation and explanations to support further research. There are a few ethical considerations for users of such technology. As with the use of fine-tuning and RAG, it is important to be aware of potential propagating biases and feedback loops. Biased data used for training and/or retrieval could lead to model outputs that reinforces such biases, impacting model quality and fairness. Although we extensively evaluated our LLM autoevaluator and found it to be safe, unbiased, and factual, users should consider whether such considerations hold true for the input source itself. Of note, even established sources such as knowledge banks may contain inaccurate or inconsistent information. Finally, though we aimed to be transparent and consistent for our data labeling tasks, using clear instructions and validations, note that data labeling tasks are inherently subject to the bias of labeler background and experiences.
Acknowledgments
@@ Appendix A Content Quality Citation Modes and the Evaluation Prompt
Here we provide the evaluation prompt template used in Figure 1, the raw autoevaluator output, and a final output example in an alternative citation mode for content quality evaluation.
@@ A.1 Content Quality Evaluation Prompt
You will be asked to providefeedback on one quality metric for a taskprompt given to another LLM and thatLLM’s output. Given the task prompt as acontext, your job is to give feedback onwhether there are any errors in the output,based on the quality metric. If thereare no errors in the output, pleasealso provide your feedback that there areno errors. Remember your main goal is toprovide feedback so that another LLM canuse your feedback to improve its generation.You need to support statements in yourfeedback that can be linked back to ANOTHERLLM’S TASK PROMPT with citations.Please make sure you read and understandthese instructions carefully. Please keepthis document open while reviewing, andrefer to it as needed.Please only generate your answer followingthe Output JSON format. Do not generateanything else.Evaluation criteria:{metric_name} ({metric_scale})- {metric_description}Respond using the following JSON format basedon the provided ANOTHER LLM’S TASK PROMPT andANOTHER LLM’S OUTPUT:Desired format:{   "answer":<answer>,   "feedback":<generated feedback>,   "statements":[   {       "statement_string":<one statement       extracted from feedback>,       "citations":[       {            "snippet":<snippet>       }]    }]}### INSTRUCTIONS:Here are the instructions to follow togenerate the response.1. Generate a feedback based on ANOTHERLLM’S TASK PROMPT and ANOTHER LLM’SOUTPUT. Explain whether the evaluationcriteria is met.    - Set "answer" to "Yes" if the evaluation    criteria is met, and "No" otherwise.    - Set "feedback" to the    generated feedback.2. For each statement in your feedback,extract relevant sentences that support thestatement from ANOTHER LLM’S TASK PROMPT.3. For each extracted sentence:    - Set "snippet" to be the extracted    sentence.4. Compile relevant snippets for thestatement:    - Set "statement_string" to be    a statement in the generated feedback.    - Set "citations" to be a list of JSON    objects containing all relevant snippets    (from step 3) supporting the    "statement_string".5. Compile the statements into a list ofJSON objects.    - Set "statements" to be a list of    JSON objects containing metadata of    all statements (from step 4) in the    feedback.[BEGIN DATA]### ANOTHER LLM’S TASK PROMPT:Summarize knowledge from transcripts afterthey’ve ended. Summarizing the key points ofthe conversation including customerissue and resolution.### Conversation:{convo}### ANOTHER LLM’S OUTPUT:{answer}[END DATA]### Response(JSON only):
@@ A.2 Content Quality Citation Modes
Using the same input example as in Figure 1, we show the raw autoevaluator output that allows users to further process according to either of the two citation modes, “post-fix with context snippet” and “inline with context snippet”, to fit their needs. With the final output in “inline with context snippet” mode already presented in Figure 1, here we show another possible final output format in “post-fix with context snippet” mode.
A.2.1 Raw Autoevaluator Output
Citations are extracted verbatim from the input context to support statements in the autoevaluator’s explanation.
"statements": [ {   "statement_string": "The answer covers all   the key points of the conversation,   including the customer’s question, the   agent’s response, and the resolution.",   "citations": ["Summarizing key points   of the conversation including customer   issue and resolution.", "Customer: Is it   possible to have my items gift wrapped?",   "Agent: We offer gift wrapping but it   cost \$4.99 per item"] }, {   "statement_string": "It also mentions the   customer’s reason for asking about gift   wrapping and the agent’s polite closing.",   "citations": ["Customer: Oh great. My   Bestie’s birthday is coming up, that’s   the reason for my questions. Thanks",   "Agent: All right. thanks for   reaching out and have a great day!"]  }]
@@ A.2.2 Final Output Example in “post-fix with context snippet” Citation Mode
The answer covers all the key points of theconversation, including the customer’s question,the agent’s response, and the resolution. It alsomentions the customer’s reason for asking aboutgift wrapping and the agent’s polite closing.[1][2][3][4][5][1]: “Summarize key points of the conversationincluding customer issue and resolution.“[2]: “Customer: Is it possible to have my itemsgift wrapped?”[3]: “Agent: We offer gift wrapping but it costs\$4.99 per item.”[4]: “Customer: Oh great. My Bestie’s birthdayis coming up, that’s the reason for my questions.Thanks.“[5]: “Agent: All right. thanks for reaching outand have a great day!” “
@@ Appendix B Details of using REC Models for General RAG citations
We demonstrate how REC models can also be used for general RAG citations in Figure 2. In the example shown, we present this question “What is photosynthesis?” to another LLM, and collect retrieved articles with ids as the context, and the generated answer from another LLM to serve as inputs. The goal is for our LLM autoevaluator to provide citations extracted verbatim from the context to support the generation from another LLM. REC models support four different citation modes, and the details can also be found in this section.
@@ B.1 RAG Citation Prompt
Your task is to providecitations for a generated responsefrom an LLM for a RAG(Retrieval Augmented Generation)application. You will be provided twosections below. The RETRIEVED CHUNKSsection is the context fromwhere you will generate citations.The LLM GENERATED ANSWER sectionis for the LLM generated answerfrom where you will generate claims.###Respond using the following JSON format:Desired format:{   "citations":[      {         "context_id":"<context_id>",         "claim": "<claims identified from         LLM GENERATED ANSWER>"      }   ]}###INSTRUCTIONSIn order to generate citations,follow these steps below:1. Review the "LLM GENERATED ANSWER"section below and extract a set of claims,VERBATIM and set the "claim" in the outputjson.2. Your task is to generate  a citationfor each of the claims identifiedin Step 1, by following the steps below:    i. For each claim, identify the    most relevant chunk and the corresponding    "context_id" from the "RETRIEVED CHUNKS"    section that supports the claim.    ii. For the identified chunk,    set the "context_id" in the output    to the corresponding context_id.    Set "context_id" to "None" if none    of the chunks in the "RETRIEVED CHUNKS"   section supports the claim.3. Ensure that you structure yourresponse to adhere to the desiredoutput JSON format.###RETRIEVED CHUNKS:{retrieved_chunks}###LLM GENERATED ANSWER:{answer}###Response (JSON only):
@@ B.3 Post-fix Citation with Context Snippet
{  "citations": [    {      "context_id": "1233",         "snippet": "Photosynthesis is the         process by which green plants and         some other organisms utilize         sunlight to synthesize         their food. This remarkable         process involves the         conversion of carbon         dioxide and water into         glucose and oxygen,         facilitated by chlorophyll"    }  ]}

### v3

@@ 6.3 General Capabilities
The RewardBench dataset assesses LLMs’ performance across several critical abilities using curated chosen-rejected response pairs (Lambert et al., 2024). The evaluation involves determining the model’s win percentage based on its ability to score the “chosen” response higher than the “rejected” one. Specifically, it’s focused on testing chat performance, safety measures, and reasoning (both in terms of code and math skills), while controlling for potential biases such as overfitting to prior datasets.
We first create prompts following the RewardBench evaluation format, where each prompt consisting of a query and two potential responses. The task for LLMs is to evaluate the two responses and determine which is better. The performance of the reward models is quantified by calculating the win percentage for chosen-rejected pairs associated with each prompt. A “win” is defined by a scenario where an LLM presents a preference for the selected answer. This quantitative measure provides a clear benchmark for evaluating model performance across tasks. By averaging the results across the 4 core sections: Chat, Chat Hard, Safety, and Reasoning, a comprehensive measure of each model’s win percentage is derived. As shown in Table 4, our REC models are able to achieve SOTA performance among all generative LLMs on the RewardBench Leaderboard, demonstrating their outstanding ability to serve as an auto-evaluator for pairwise comparisons.
@@
This is a benchmark for measuring the grounding capabilities of auto-evaluators. Given a reference document and a claim, the auto-evaluator determines if the claim is fully supported by the document. This holistic benchmark combines 10 attribution datasets used in recent studies on LLM factuality. Given some of the models such as GPT-3.5 have smaller context windows, we removed examples longer than 16K tokens from our evaluation. We use the same prompt instructions across categories, as shown in Appendix G. To reduce model API costs, we randomly sampled 256 examples per evaluation task similar to the approach used in the FLAMe (Vu et al., 2024b).
Table 5 presents our attribution results on LLM-AggreFact (Tang et al., 2024b), categorized into four common use-cases: (1) LLM-FactVerify: fact verification of LLM-generated responses, (2) Wiki-FactVerify: evaluating correctness of Wikipedia claims, (3) Summarization: assessing faithfulness of summaries, and (4) Long-form QA: evaluating long-form answers to questions. REC-12B and REC-70B outperform all other models in all four categories, and REC-8B outperform all the other similar-sized models (Mistral-7B and Mistral-Nemo), demonstrating their strong performance in attribution on diverse datasets.
@@ 6.4 Bias Testing
Recent studies have found that LLM-as-a-Judge often exhibits cognitive biases, such as preferences for verbosity, egocentrism, bandwagon, and an overly authoritative tone (Wang et al., 2024a; Koo et al., 2024; Zheng et al., 2024; Chen et al., 2024). To investigate the biases of the compared models, we evaluate them on the CoBBLEr benchmark (Cognitive Bias Benchmark for LLMs as EvaluatoRs) (Koo et al., 2024). This dataset is designed to evaluate the quality and reliability of LLMs when used as automated evaluators in a question-answering (QA) setting. It assesses the presence of six cognitive biases, both implicit and induced, when LLMs are tasked with ranking responses generated by various other models. CoBBLEr’s core objective is to identify the extent of bias in LLM evaluation outputs.
The CoBBLEr dataset includes 50 QA instructions, randomly selected from two well-established benchmarks: BIG-bench (bench authors, 2023) and ELI5 (Fan et al., 2019). 16 LLMs, both open and closed-source models, generate responses to these instructions. The evaluations involve pairwise comparisons between the responses of two models, wherein each model also acts as an evaluator to rank its own and others’ outputs. The biases tested are categorized into two groups: (1) Implicit biases, such as egocentric bias (where a model tends to prefer its own outputs), and (2) Induced biases, such as order bias, where the ranking of responses is influenced by their order in the evaluation.
@@ 7 Conclusions
LLMs excel in generating coherent and high-quality text for various tasks, but evaluating content quality is challenging due to issues like factual inaccuracies and hallucination. This paper introduces three fine-tuned models—REC-8B, REC-12B, and REC-70B—designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, completeness. These models not only provide ratings for these metrics but also offer detailed explanations and verifiable citation, i.e., Ratings, Explanations, and citation (REC), thereby enhancing trust in the content. To achieve this goal, a curated dataset for explanations and citation fine-tuning is proposed to facilitate future research, which is the first public dataset containing both content quality citation and RAG citation. The models support various citation modes, including Post-fix citation, Post-fix citation mode with snippet, Inline citation, and Inline citation mode with snippet to balance the trade-off between latency and granularity of citation. Extensive evaluations on diverse benchmarks (i.e., ALCE, ExpertQA, ABCD summarization, RewardBench, LLM-AggreFact, and CoBBLEr) demonstrate that the fine-tuned LLM auto-evaluator, REC-70B, outperforms state-of-the-art models. It excels in content evaluation, delivering clear explanations and ensuring citation accuracy.
Acknowledgements
@@
Our work introduces the innovative Rate, Explain, and Cite (REC) model for trusted automated AI evaluations, offering significant benefits but with a few limitations that future research will address. First, our work is limited to the baselines we compare, as non-LLM citation methods, for instance, approaches that leverage sentence embedding (Reimers & Gurevych, 2019) similarity to identify and cite the most semantically relevant sources were not included. Including such techniques could serve as valuable baselines to further validate our model’s generalization and effectiveness. Second, the need for post-generation explanations and citation may introduce latency. Future work will compare latencies across models and explore optimizations to improve speed. We are also considering fine-tuning the snippet output using starting and ending character numbers to optimize for latency. Finally, even though the 12B base model is capable of generating multi-lingual output, our current evaluations are limited to English datasets. Expanding to multilingual evaluations is a key goal for future research.
Impact Statement
Our contributions present a novel general-purpose auto-evaluator that supports various modes of citation, as well as a dataset with citation and explanations to support further research. There are a few ethical considerations for users of such technology. As with the use of fine-tuning and RAG, it is important to be aware of potential propagating biases and feedback loops. Biased data used for training and/or retrieval could lead to model outputs that reinforces such biases, impacting model quality and fairness. Although we extensively evaluated our LLM auto-evaluator and found it to be safe, unbiased, and factual, users should consider whether such considerations hold true for the input source itself. Of note, even established sources such as knowledge banks may contain inaccurate or inconsistent information. Finally, though we aimed to be transparent and consistent for our data labeling tasks, using clear instructions and validations, note that data labeling tasks are inherently subject to the bias of labeler background and experiences.
@@ Appendix A Content Quality Citation Modes and the Evaluation Prompt
Here we provide the evaluation prompt template used in Figure 1, the raw auto-evaluator output, and a final output example in an alternative citation mode for content quality evaluation.
@@ A.1 Content Quality Evaluation Prompt
You will be asked to provide
feedback on one quality metric for a task
prompt given to another LLM and that
LLM’s output. Given the task prompt as a
context, your job is to give feedback on
whether there are any errors in the output,
based on the quality metric. If there
are no errors in the output, please
also provide your feedback that there are
no errors. Remember your main goal is to
provide feedback so that another LLM can
use your feedback to improve its
generation.
You need to support statements in your
feedback that can be linked back to ANOTHER
LLM’S TASK PROMPT with citations.
Please make sure you read and understand
these instructions carefully. Please keep
this document open while reviewing, and
refer to it as needed.
Please only generate your answer following
the Output JSON format. Do not generate
anything else.
Evaluation criteria:
{metric_name} ({metric_scale})
- {metric_description}
Respond using the following JSON format
based on the provided ANOTHER LLM’S
TASK PROMPT and ANOTHER LLM’S OUTPUT:
Desired format:
   "answer":<answer>,
   "feedback":<generated feedback>,
   "statements":[
       "statement_string":<one statement
       extracted from feedback>,
       "citations":[
            "snippet":<snippet>
### INSTRUCTIONS:
Here are the instructions to follow to
generate the response.
1. Generate a feedback based on ANOTHER
LLM’S TASK PROMPT and ANOTHER LLM’S
OUTPUT. Explain whether the evaluation
criteria is met.
    - Set "answer" to "Yes" if the
    evaluation criteria is met,
    and "No" otherwise.
    - Set "feedback" to the
    generated feedback.
2. For each statement in your feedback,
extract relevant sentences that support
the statement from ANOTHER LLM’S
TASK PROMPT.
3. For each extracted sentence:
    - Set "snippet" to be the extracted
    sentence.
4. Compile relevant snippets for the
statement:
    - Set "statement_string" to be
    a statement in the generated feedback.
    - Set "citations" to be a list of JSON
    objects containing all relevant
    snippets (from step 3) supporting the
    "statement_string".
5. Compile the statements into a list of
JSON objects.
    - Set "statements" to be a list of
    JSON objects containing metadata of
    all statements (from step 4) in the
    feedback.
[BEGIN DATA]
### ANOTHER LLM’S TASK PROMPT:
Summarize knowledge from transcripts
after they’ve ended. Summarizing the key
points of the conversation including
customer issue and resolution.
### Conversation:
{convo}
### ANOTHER LLM’S OUTPUT:
{answer}
[END DATA]
### Response(JSON only):
@@ A.2 Content Quality Citation Modes
Using the same input example as in Figure 1, we show the raw auto-evaluator output that allows users to further process according to either of the two citation modes, “post-fix with context snippet” and “inline with context snippet”, to fit their needs. With the final output in “inline with context snippet” mode already presented in Figure 1, here we show another possible final output format in “post-fix with context snippet” mode.
A.2.1 Raw auto-evaluator Output
Citations are extracted verbatim from the input context to support statements in the auto-evaluator’s explanation.
"statements": [
   "statement_string": "The answer covers
   all the key points of the conversation,
   including the customer’s question, the
   agent’s response, and the resolution.",
   "citations": ["Summarizing key points
   of the conversation including customer
   issue and resolution.", "Customer: Is it
   possible to have my items gift wrapped?"
   ,"Agent: We offer gift wrapping but it
   cost \$4.99 per item"]
   "statement_string": "It also mentions the
   customer’s reason for asking about gift
   wrapping and the agent’s polite closing."
   ,"citations": ["Customer: Oh great. My
   Bestie’s birthday is coming up, that’s
   the reason for my questions. Thanks",
   "Agent: All right. thanks for
   reaching out and have a great day!"]
@@ A.2.2 Final Output Example in “post-fix with context snippet” Citation Mode
The answer covers all the key points of the
conversation, including the customers
question, the agents response, and the
resolution. It also mentions the customers
reason for asking about gift wrapping and
the agents polite closing.
[1][2][3][4][5]
[1]: "Summarize key points of the
conversation including customer issue
and resolution."
[2]: "Customer: Is it possible to have
my items gift wrapped?"
[3]: "Agent: We offer gift wrapping but
it costs \$4.99 per item."
[4]: "Customer: Oh great. My Besties
birthday is coming up, thats the reason
for my questions. Thanks."
[5]: "Agent: All right. thanks for
reaching out and have a great day!"
@@ Appendix B Details of using REC Models for General RAG citations
We demonstrate how REC models can also be used for general RAG citations in Figure 2. In the example shown, we present this question “What is photosynthesis?” to another LLM, and collect retrieved articles with ids as the context, and the generated answer from another LLM to serve as inputs. The goal is for our LLM auto-evaluator to provide citations extracted verbatim from the context to support the generation from another LLM. REC models support four different citation modes, and the details can also be found in this section.
@@ B.1 RAG Citation Prompt
Your task is to provide
citations for a generated response
from an LLM for a RAG
(Retrieval Augmented Generation)
application. You will be provided two
sections below. The RETRIEVED CHUNKS
section is the context from
where you will generate citations.
The LLM GENERATED ANSWER section
is for the LLM generated answer
from where you will generate claims.
Respond using the following JSON format:
Desired format:
   "citations":[
         "context_id":"<context_id>",
         "claim": "<claims identified from
         LLM GENERATED ANSWER>"
###INSTRUCTIONS
In order to generate citations,
follow these steps below:
1. Review the "LLM GENERATED ANSWER"
section below and extract a set of claims,
VERBATIM and set the "claim" in the output
json.
2. Your task is to generate  a citation
for each of the claims identified
in Step 1, by following the steps below:
    i. For each claim, identify the
    most relevant chunk and the
    corresponding "context_id" from
    the "RETRIEVED CHUNKS" section that
    supports the claim.
    ii. For the identified chunk,
    set the "context_id" in the output
    to the corresponding context_id.
    Set "context_id" to "None" if none
    of the chunks in the
    "RETRIEVED CHUNKS" section supports
    the claim.
3. Ensure that you structure your
response to adhere to the desired
output JSON format.
RETRIEVED CHUNKS:
{retrieved_chunks}
LLM GENERATED ANSWER:
{answer}
Response (JSON only):
@@ B.3 Post-fix Citation with Context Snippet
  "citations": [
      "context_id": "1233",
         "snippet": "Photosynthesis is
         the process by which green
         plants and some other organisms
         utilize sunlight to synthesize
         their food. This remarkable
         process involves the
         conversion of carbon
         dioxide and water into
         glucose and oxygen,
         facilitated by chlorophyll"